OpenStudy (anonymous):

elimination method 4x+y=-4 6x+y=-10

OpenStudy (anonymous):

y = -4 -4x plug it into 6x+y=-10 6x -4 -4x = -10 2x = -6 x = -3 return to y = -4 -4x and get : y = -4 + 12 = 8 (-3,8)

OpenStudy (anonymous):

that is the substitution method

OpenStudy (anonymous):

Solve for one, either X or Y, then plug into the other equation. So if we solve for y in the first equation, we have 4x+y=-4 or y = -4 - 4x Now we can plug -4 + 4x in eveywhere we see Y in the second equation We have 6x+y=-10 6x + (-4 - 4x) = -10 This gives us 6x -4 - 4x = -10 Which is 2x -4 = -10 2x = -6 x = -3 And we can take this x and plug it back into our answer for y, which was y = -4 - 4x giving us y = -4 - 4(-3) so we have y = -4 +12 and so y=8 So we have x=-3, and y=8

OpenStudy (anonymous):

oh elimination so : 4x+y=-4 6x+y=-10 subtract the second from the first to get : -2x = 6 x= -3 then plug it into the first -12 +y = -4 y = 8

OpenStudy (anonymous):

to eliminate, say, the y-variables, mutliply the first equation by -1 and then form a linear combination of the two: -4x - y = 4 6x + y =-10 ========= 2x =-6 x=-3 now you can sub x=-3 into any eqaution above that has a y-variable in it; i choose the second given eqation: 6(-3) + y = -10 y=-10+18 y=8 the solution is x=-3; y=-8; or as an ordered pair (-3, 8)